Albin has quasi linear preferences and he loves pretzels. His inverse demand function for pretzels is p(x) = 49 − 6x, where x is the number of pretzels that he consumes. He is currently consuming 8 pretzels at a price of £1 per pretzel. What is the change in Albin’s consumer surplus if the price of pretzels rises to £7 per pretzel?
1
Expert's answer
2017-11-22T14:57:06-0500
(0.5*(49-1)*8)-(0.5*(49-7)*7)=192-147=45